0 0.051 6.4 0.164 12.9 0.28 19.3 0.382 25.7 0.499 Graph Absorbance (y-axis) vs Concentration (x-axis)....

0 0.051
6.4 0.164
12.9 0.28
19.3 0.382
25.7 0.499

Graph Absorbance (y-axis) vs Concentration (x-axis).

What is the slope of the line?

What is the concentration of the solution if the solution is 0.317?

what would be the Absorptivity (ɛ) of the substance if the path length of the cuvette, b = 1 cm?
Answer #1

Answer:

concentration absorbance 0 0.051 6.4 0.164 12.9 0.28 19.3 0.382 25.7 0.499

Plot is,

y = 0.0173x + 0.0524 R= 0.9996 Absorbance 0.3 5 10 15 20 concentration (micromolar) 25 30

Answer 2:

Slope of the line is = 0.0173

(unit is per micromolar or \mu M-1)

Answer 3:

If absorbance = 0.317 = y

equation of ine is,       y = 0.0173 x      +    0.0524

so x = (y - 0.0524) / 0.0173 = (0.317 - 0.0524) / 0.0173 = 15.3

If absorbance is 0.317, then concentration of solution will be 15.3 \mu M.

Answer 4:

if path length = 1 cm

then molar absorptivitybity = 0.0173 cm-1 \mu M-1   = 0.0173 cm-1 x 106 M-1 = 17300 cm-1 M-1

Molar absorptivity = 17300 cm-1 M-1